OpenStudy (anonymous):

Write the equation of a hyperbola with vertices (-3, 0) and (3, 0) and co-vertices (0, -5) and (0, 5).

OpenStudy (nurali):

Given hyperbola has a horizontal transverse axis. Its standard form: (x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of centre For given hyperbola: centre:(0,0) length of horizontal transverse axis=6=2a a=3 a^2=9 length of co-vertices or conjugate axis=10=2b b=5 b^2=25 Equation: x^2/9-y^2/25=1

OpenStudy (anonymous):

thnks so much... i been over here crying and struggling because i an taking an exam online and i have to pass it.. i was wondering if you can help me out

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

ok im gonna give you the answers so you can help me

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

huh?